A balloon of carbon dioxide gas has a volume of 14.1 L and contains 5.3 moles of carbon dioxide. A leak develops some carbon dioxide escapes, leaving only 2.2 moles of carbon dioxide inside the balloon. What is the new volume of the balloon? Enter only the numeric value for your answer (no units).
Answer #1

Given:

Vi = 14.1 L

ni = 5.3 mol

nf = 2.2 mol

use:

Vi/ni = Vf/nf

14.1 L / 5.3 mol = Vf / 2.2 mol

Vf = 5.85 L

Answer: 5.85 L